### v2

Abstract
This paper presents a theoretical framework that explains why fine-tuning small, randomly selected subnetworks (slices) within pre-trained models is sufficient for downstream adaptation. We prove that pretrained networks exhibit a universal winning slice property, arising from two phenomena: (1) spectral balance—the eigenspectra of different weight matrix slices are remarkably similar—and (2) high task energy—their backbone representations (pretrained weights) retain rich, task-relevant features. This leads to the Universal Winning Slice Hypothesis, which provides a theoretical foundation for parameter-efficient fine-tuning (PEFT) in large-scale models. Inspired by this, we propose SliceFine, a PEFT method that uses this inherent redundancy by updating only selected slices of the original weights—introducing zero new parameters, unlike adapter-based approaches. Empirically, SliceFine matches the performance of SOTA PEFT methods across various language and vision tasks, while significantly improving training speed, memory efficiency, and model compactness. Our work bridges theory and practice, offering a theoretically grounded alternative to existing PEFT techniques.
Figure 1: (Left) Winning Tickets. In a pretrained network, a randomly chosen slice of a layer $W^{(\ell)}\in\mathbb{R}^{d_{\ell}\times d_{\ell-1}}$ acts as a local winning ticket: tuning only that slice lowers the loss while keeping the backbone frozen. A few such slices (row, column, or row-column) selected across layers constitute a global winning ticket. (Right) SliceFine. At step $t$, only a slice of the weight matrix $W^{(\ell)}$ is updated; all other entries remain fixed. Every NN steps, we activate a new slice at a different position for learning; the previously active slice retains its learned update but is frozen. Top: column sweep—the slice slides across columns. Bottom: row–column alternation—the slice alternates between a column block and a row block to cover complementary directions. Similarly, in row sweep—the slice slides across rows. This schedule updates only a tiny portion of the model at a time while gradually covering many regions; applying it across several layers yields a global winner.
1 Introduction
Large pretrained models work well across many tasks, yet we still lack a simple picture of why small changes are often enough to adapt them. Popular parameter-efficient fine-tuning (PEFT) methods add small modules (e.g., adapters, low-rank layers) and train only those parts while freezing the rest (Hu et al., 2021; Kowsher et al., 2024b; Zhang et al., 2023b; Liu et al., 2024b; Prottasha et al., 2025). This raises a basic question: why do these small modules work, and do we actually need to add new parameters—or is there already enough useful structure inside the pretrained weights themselves?
This paper addresses this question by introducing the Universal Winning Slice Hypothesis, which provides a simple, testable explanation for fine-tuning behavior in large-scale pretrained models and guides a very light form of adaptation. First, we briefly fix terminology and state the Universal Winning–Slice Hypothesis.
Terminology. Let $f_{\theta_{0}}:\mathbb{R}^{d}\to\mathbb{R}^{k}$ denote a pretrained neural network of depth $L$, where each layer $\ell$ is parameterized by a weight matrix $W^{(\ell)}\in\mathbb{R}^{d_{\ell}\times d_{\ell-1}}$. A slice of a layer $l$ is a contiguous set of rows or columns of a single weight matrix $W^{(\ell)}$, specified by a mask $M^{(\ell)}\in\{0,1\}^{d_{\ell}\times d_{\ell-1}}$. A winning slice (local winner) is a slice whose update alone reduces the loss at $\theta_{0}$. A slice set $\mathcal{T}=\{M_{i}\}_{i=1}^{m}$ is a small collection of slices across layers. When the joint update over $\mathcal{T}$ attains near full fine-tuning performance, $\mathcal{T}$ is a global winning ticket. We call $r$ the slice rank (number of selected columns or rows). Slice rank is not the matrix rank of $W^{(\ell)}$; it measures the slice width (capacity).
Universal Winning–Slice Hypothesis (UWSH). In a dense, pretrained network, any random slice with sufficient width has a local winning ticket: training only that slice while freezing the rest improves downstream performance (a local winner). Moreover, tuning a small set of such slices across layers can match full fine-tuning accuracy while updating far fewer parameters (a global winner).
This view differs from the Lottery Ticket Hypothesis (LTH) (Frankle & Carbin, 2019; Yu et al., 2021; Chen et al., 2020), which posits a special, sparse subnetwork (“winning ticket”) that must be found by pruning and then trained—often from its original initialization—to match full–network performance. Most analyses of LTH rely on carefully identifying such subnetworks via iterative pruning by randomly (Bai et al., 2022) or importance-based selection (You et al., 2022). In contrast, we study large-scale, dense, pretrained networks and establish a theoretical framework showing that a random simple slice can serve as a winning ticket. Consequently, our goal is to fine-tune only these slices—without adding new parameters—as a PEFT strategy for adapting pretrained models to downstream tasks. The reason every sufficiently wide slice can be a winning ticket is twofold: First, spectral balance. Take any layer and split its weight matrix into simple groups of rows or columns (slices). In Figure 2, if we look at each group’s covariance and its eigenvalues, we find that the average size of the eigenvalues and the way they decay are very similar across groups, implying that all slices have comparable capacity for fine-tuning. Second, high task energy. After pretraining, the backbone features of a model already line up a large share of their variation with directions that are useful for downstream tasks. This is reflected in the high cumulative explained variance of the top PCA components of the centered features (details in Lemma 2.5), or, equivalently, a “lazy” NTK spectrum where a few top directions dominate. When most of the useful signal lives in a small set of directions, a small slice that has enough rank will inevitably touch those directions and can move the model in the right way.
Together, these two facts give a simple rule of thumb. Because slices are balanced in strength, which slice we pick is not critical. Because the backbone already concentrates task energy, a slice in $W^{(\ell)}$ overlaps with the task-relevant subspace, produces a nonzero restricted gradient, and reduces the loss. This picture also clarifies common PEFT observations: adapters help because the backbone already carries most of what is needed (see Corollary D.2).
As every sufficiently wide slice in $W^{(\ell)}$ has a local winning ticket, can we fine-tune only a slice? Based on the UWSH, we propose SliceFine, which trains a set of slices $\mathcal{T}$ across different layers (to get a global winning ticket) while keeping the backbone frozen and adds no new parameters. The per-layer cost scales with the slice rank, and in practice very small ranks—often r=1r{=}1—already match full fine-tuning or strong PEFT baselines on downstream tasks. Although Figure 12 shows good accuracy on downstream tasks when training only a fixed slice in $W^{(\ell)}$, to cover more directions over time we freeze the active slice after a fixed number of steps $N$ and then activate a new slice at a different position, effectively performing block-coordinate descent across the layer. In practice, we do not need to traverse all possible positions; switching the active slice every $N=500$ training steps and making only 551010 switches in total achieves optimal performance across all of our experiments. Figure 1 (right top) illustrates a column-slice schedule: at t=0t{=}0 the first column group is trained; after $N$ steps the slice shifts to the next column group, and so on. Figure 1 (right bottom) shows an alternating schedule: at t=0t{=}0 a column slice is trained; at t=Nt{=}N the slice switches to a row group; subsequent shifts continue alternating between columns and rows.
Our contributions are summarized as follows: (i) A theoretical and testable hypothesis, UWSH: in pretrained networks, an arbitrary sufficiently wide slice is a local winning ticket; multiple slices across layers form a global winning ticket. (ii) A PEFT method SliceFine by following UWSH: trains only a set of slice $\mathcal{T}$ across different layers, with no new parameters. (iii) Broad empirical study across language and vision showing accuracy on par with—or better than—strong PEFT baselines, with lower memory, faster throughput, and smaller model footprints. (iv) Ablations that connect rank choice, slice selection, switching interval, initialization, and backbone pruning to the simple picture above, using explained variance/NTK laziness as a guide.
2 Winner Slices in Pretrained Networks
Let $Y\in\mathbb{R}^{c\times n}$ be the label encoding matrix for $n$ downstream examples across $c$ classes, and let $P_{Y}\in\mathbb{R}^{n\times n}$ be the orthogonal projector onto the column space of $Y$. The task covariance at layer $\ell$ is $\Sigma_{\mathrm{task}}:=\frac{1}{n}\,\Phi_{\ell}\,P_{Y}\,\Phi_{\ell}^{\top}$ and task subspace $k_{\mathrm{task}}:=\mathrm{rank}(\Sigma_{\mathrm{task}}).$
Consider a network $f_{\theta_{0}}$ is pretrained on a large dataset so that, for each layer $\ell$, the intermediate representations $\phi_{\ell}(x)\in\mathbb{R}^{d_{\ell}}$ satisfy: $\mathrm{rank}(\Phi_{\ell})=\mathrm{rank}([\phi_{\ell}(x_{1}),\dots,\phi_{\ell}(x_{n})])\geq k_{\mathrm{task}},$ where $\Phi_{\ell}\in\mathbb{R}^{d_{\ell}\times n}$ is the matrix of downstream representations and $k_{\mathrm{task}}$ is the intrinsic dimension of the downstream task in the representation space at layer $\ell$, i.e., the smallest number of orthogonal directions needed to achieve perfect linear classification of the labels.
Formally, the slice parameters $M^{(\ell)}\odot W^{(\ell)}$ (where $\odot$ denotes element-wise multiplication) are trainable, while all other entries remain fixed at their pretrained values. Unlike pruning in LTH (Yu et al., 2021; Chen et al., 2020), the frozen weights are not zeroed out; they continue to contribute during forward propagation, providing a representational scaffold that allows the trainable slice to adapt effectively. The learned increment is $\Delta W^{(\ell)}=M^{(\ell)}\odot U^{(\ell)},$ where $U^{(\ell)}$ denotes the gradient-driven update. Given a downstream dataset $D=\{(x_{i},y_{i})\}_{i=1}^{n}$ and convex per-sample loss $\ell$, the fine-tuning objective is: $\mathcal{L}(\theta)=\frac{1}{n}\sum_{i=1}^{n}\ell(f_{\theta}(x_{i}),y_{i}).$
Let $U_{k_{\mathrm{task}}}\in\mathbb{R}^{d_{\ell}\times k_{\mathrm{task}}}$ denote the top $k_{\mathrm{task}}$ left singular vectors of $\Phi_{\ell}$, spanning the task-relevant subspace. The projection operator is $P_{U_{k_{\mathrm{task}}}}=U_{k_{\mathrm{task}}}U_{k_{\mathrm{task}}}^{\top}.$ The quantity $k_{\mathrm{task}}$ captures the minimum number of feature-space directions required for perfect linear separation of the downstream labels. In SliceFine, we must determine whether each slice of a pretrained weight matrix overlaps enough with the $k_{\mathrm{task}}$-dimensional task subspace to reduce downstream loss. The lemma below shows that, in pretrained networks $f_{\theta_{0}}$, slices from the $W^{(\ell)}$ have similar average spectral energy and comparable eigenvalue decay—a property we call spectral balance.
Lemma 2.1 (Spectral Balance Across Slices).
Consider a pretrained layer $W^{(\ell)}$ partitioned into $k$ disjoint groups $\{W_{g}\}_{g=1}^{k}$. Let $\Sigma_{g}:=W_{g}W_{g}^{\top}$ and let $\lambda_{1}(\Sigma_{g})\geq\dots\geq\lambda_{d_{\ell}/k}(\Sigma_{g})$ denote its eigenvalues in descending order. Although each group is anisotropic—its eigenvalues decay from large to small—the groups exhibit spectral balance in the sense that: $\frac{\frac{1}{d_{\ell}/k}\sum_{i=1}^{d_{\ell}/k}\lambda_{i}(\Sigma_{g})}{\frac{1}{d_{\ell}/k}\sum_{i=1}^{d_{\ell}/k}\lambda_{i}(\Sigma_{g^{\prime}})}\approx 1,\quad\forall g,g^{\prime}\in\{1,\dots,k\},$ and their decay profiles are boundedly similar: $\max_{1\leq i\leq d_{\ell}/k}\frac{\left|\lambda_{i}(\Sigma_{g})-\lambda_{i}(\Sigma_{g^{\prime}})\right|}{\lambda_{i}(\Sigma_{g^{\prime}})}\leq\rho,$ for some small $\rho\ll 1$ independent of $g,g^{\prime}$. Thus, no group is disproportionately “weak” or “strong” in average spectral energy, and each slice retains non-trivial overlap with the task-relevant subspace.
This property guarantees that no slice is degenerate with respect to the task subspace: all slices retain comparable spectral energy and a non-zero projection onto task-relevant directions. Consequently, each slice admits a non-zero restricted gradient with respect to the downstream loss and can independently reduce the loss—i.e., it qualifies as a local winning slice.
Figure 2: Eigenvalue spectra of FFN, Key, Query, and Value weight matrices from different layers of a pretrained RoBERTa-base model. For each matrix, weights are partitioned into groups, and the eigenvalues of the within-group covariance $\Sigma_{g}=W_{g}^{(\ell)}W_{g}^{(\ell)\top}$ are plotted in descending order.
Figure 2 shows that, although each group’s eigen-spectrum is anisotropic—dominated by a few large eigenvalues—the decay profiles and the average spectral energy are nearly identical across groups. Groups are constructed by partitioning the weight matrix into $n$ equal blocks: for a row-wise partition (first two figures) we split the $d_{\ell}$ rows into $n$ groups (each of rank $r=d_{\ell}/n$), and for a column-wise partition (last two figures) we split the $d_{\ell-1}$ columns into $n$ groups (each of rank $r=d_{\ell-1}/n$). Empirically, the average inter-group variance metric $\rho$ computed across all groups and eigen-components is $0.00125$, $0.00192$, $0.00255$, and $0.00102$, respectively, indicating negligible dispersion. This supports the spectral balance property of Lemma 2.1, implying that no group is disproportionately weak or strong.
Spectral balance is the key precondition that motivates the following two definitions. We first define a local winning slice to capture the ability of a slice to reduce downstream loss, and then a global winning ticket to describe a set of slices $\tau$ across different layers whose joint optimization achieves full fine-tuning performance.
Definition 2.2 (Local Winner).
A mask $M$ is a local winning slice if there exists an update $U$ such that $\mathcal{L}(\theta_{0}+M\odot U)\leq\mathcal{L}(\theta_{0})-\delta$ for some $\delta>0$, and the restricted gradient is non-zero: $\|\nabla_{M}\mathcal{L}(\theta_{0})\|_{2}=\Bigg\|\frac{\partial\mathcal{L}(\theta)}{\partial(M\odot W^{(\ell)})}\Bigg|_{\theta=\theta_{0}}\Bigg\|_{2}>0.$ Under spectral balance, every slice has comparable spectral energy and retains overlap with the task subspace, ensuring its restricted gradient is non-zero and enabling it to independently reduce loss.
Definition 2.3 (Global Winner).
A set of masks $\mathcal{T}=\{M_{i}\}_{i=1}^{m}$ forms a global winning ticket if there exist updates $\{U_{i}\}_{i=1}^{m}$ such that $\mathcal{L}\Big(\theta_{0}+\sum_{i=1}^{m}M_{i}\odot U_{i}\Big)\leq\epsilon,$ for some $\epsilon>0$, achieving performance close to full fine-tuning while updating only a small fraction of weights.
While the local winning slice definition ensures each slice can make independent progress, the global winning ticket definition captures the compositional effect: if different slices contribute complementary directions in the task subspace, their combined span can cover the full $k_{\mathrm{task}}$ subspace, matching full fine-tuning performance.
Theorem 2.4 (Universal Winning Ticket).
Let $f_{\theta_{0}}$ be a pretrained network of depth $L$. For each layer $\ell$, let $\Phi_{\ell}=[\phi_{\ell}(x_{1}),\dots,\phi_{\ell}(x_{n})]$ have singular values $\sigma_{1}\geq\sigma_{2}\geq\dots$ satisfying $\frac{\sum_{j=1}^{k_{\mathrm{task}}}\sigma_{j}^{2}}{\sum_{j}\sigma_{j}^{2}}\geq\eta,$ for some $\eta\in(0,1)$. Assume further that every slice vector $w_{j}$ of $W^{(\ell)}$ has non-trivial projection into the top-$k_{\mathrm{task}}$ subspace: $\|P_{U_{k_{\mathrm{task}}}}w_{j}\|_{2}^{2}\geq\gamma>0.$
Then: 1. (Local winners) For any binary mask $M$ selecting a slice of weights in layer $\ell$, if the restricted Jacobian $J_{M}(x):=\nabla_{M\odot W^{(\ell)}}f_{\theta_{0}}(x)$ satisfies $\mathrm{rank}(J_{M})\geq k_{\mathrm{task}}>0$, then the gradient on the slice is nonzero: $\left\|\frac{\partial\mathcal{L}(\theta)}{\partial(M\odot W^{(\ell)})}\right\|_{F}\Bigg|_{\theta=\theta_{0}}>0.$ Consequently, there exists a perturbation $U$ supported on $M$ such that, for sufficiently small $\eta>0$, $\mathcal{L}(\theta_{0}+\eta\,M\odot U)<\mathcal{L}(\theta_{0})-\delta,$ making the slice a local winner. This follows from spectral balance and projection properties, which ensure $J_{M}$ retains task-relevant directions.
2. (Global tickets) There exists a collection $\{M_{i}\}_{i=1}^{m}$, with $m\ll\sum_{\ell}d_{\ell}$, such that joint or sequential optimization yields $\mathcal{L}\Big(\theta_{0}+\sum_{i=1}^{m}M_{i}\odot U_{i}\Big)\leq\epsilon,$ for arbitrarily small $\epsilon$. This holds because the combined Jacobians $\dim\mathrm{span}(J_{M_{1}}(x),\dots,J_{M_{m}}(x))\to k_{\mathrm{task}},$ incrementally span the task-relevant subspace, enabling near full fine-tuning performance while updating only a small subset of weights.
Theorem 2.4 follows naturally: spectral balance ensures that any slice is a local winner (Part 1), while the finite rank $k_{\mathrm{task}}$ determines the number of slices required to assemble a global ticket (Part 2). Details of the proof are provided in Appendix A.
We now examine how large a slice must be to act as a winner. Empirically, our ablation study (Appendix E.1) shows that training a slice with rank one ($r=1$ i.e., updating a single column or a single row of a pretrained weight matrix) is often sufficient. Theoretically, we find that this rank requirement is determined by the spectral structure of the representation space, which can be understood via the relationship between the linearized NTK kernel and the PCA decomposition of the features. We observe a strong dependence on the domain. When the feature spectrum exhibits high cumulative explained variance (CEV) and the model remains in a lazy NTK regime after fine-tuning, a slice rank $r$ suffices to capture the task directions. Conversely, when the spectrum has lower CEV and the NTK is less lazy, larger slices are needed to cover enough task-relevant directions. Since, in the linearized regime, the NTK spectrum matches the PCA spectrum of the centered features (Lemma 2.5), the PCA curve gives a direct guide: a steeper curve suggests a slice suffices; a flatter curve calls for a larger slice.
Lemma 2.5 (PCA Decomposition of the Representation/Linearized NTK Kernel).
Let $H\in\mathbb{R}^{n\times d}$ be hidden representations and $\tilde{H}:=H-\frac{1}{n}\mathbf{1}\mathbf{1}^{\top}H$ the centered features. Let $\Sigma:=\frac{1}{n-1}\tilde{H}^{\top}\tilde{H}$ with eigendecomposition $\Sigma=V\Lambda V^{\top}$, where $V\in\mathbb{R}^{d\times r}$ has orthonormal columns and $\Lambda=\mathrm{diag}(\lambda_{1},\dots,\lambda_{r})$ with $\lambda_{i}>0$.
Define $P:=\tilde{H}V\in\mathbb{R}^{n\times r}$. Then the (centered) feature Gram matrix (a.k.a. linearized representation kernel) $K:=\tilde{H}\tilde{H}^{\top}\in\mathbb{R}^{n\times n}$ admits the exact decomposition $\boxed{K\;=\;PP^{\top}}.$
Moreover, the nonzero eigenvalues of $K$ are $(n-1)\lambda_{1},\dots,(n-1)\lambda_{r}$. Consequently, for any $k\leq r$$\frac{\sum_{i=1}^{k}\lambda_{i}}{\sum_{j=1}^{r}\lambda_{j}}\;=\;\frac{\sum_{i=1}^{k}\mu_{i}}{\sum_{j=1}^{r}\mu_{j}},\qquad\mu_{i}:=(n-1)\lambda_{i},$ so PCA explained-variance ratios in feature space match those of $K$.
Corollary 2.6 (Minimal Slice Rank from PCA/NTK Spectrum).
Let $k_{\mathrm{task}}(\tau)$ be the smallest integer $m$ such that $\sum_{i=1}^{m}\lambda_{i}/\sum_{j=1}^{r}\lambda_{j}\geq\tau$ for a threshold $\tau\in(0,1]$. If the slice rank $r_{\mathrm{slice}}\geq k_{\mathrm{task}}(\tau)$, then with high probability that the slice intersects the task-relevant subspace, and thus can serve as a local winner. For familiar domains, $k_{\mathrm{task}}(\tau)$ is small due to a steep PCA/NTK spectrum; for unfamiliar domains, it is larger due to a flatter spectrum.
The proof of Lemma 2.5 is given in Appendix B. Additional discussion and empirical evidence connecting slice rank, NTK laziness, and PCA variance profiles are provided in Appendix C.
Now we turn our attention to analyzing how much task energy the frozen backbone already carries for an $r$-rank slice to be a winner. We measure this with the PCA/NTK cumulative explained variance of the features: when the CEV is high, most of the feature energy lies in a small set of directions that separate the labels (the “task subspace”). From a small calibration set, estimate $k_{\mathrm{task}}(\tau)$: the fewest principal components that explain at least a fraction $\tau$ of the feature variance. Then set the slice size $r_{\text{slice}}\geq k_{\mathrm{task}}(\tau)$. Under the same spectral-balance assumption as in Lemma D.1, any slice of this size has nonzero overlap with the task subspace, so its gradient is nonzero and the loss decreases when we update it. In short: if the backbone shows high task energy, a slice that is large enough will train well. This also explains the failure modes. If we weaken the backbone (e.g., heavy pruning or strong domain shift), the CEV curve becomes flatter, the effective task subspace grows, and the slice’s overlap shrinks. Lemma D.1 then predicts smaller gradients and smaller guaranteed improvement. To recover performance we must increase $r_{\text{slice}}$ or combine several slices until their total span reaches $k_{\mathrm{task}}(\tau)$. These trends match our experiments: familiar domains (high CEV) work with very small slices, while unfamiliar or degraded backbones require larger slices (Details in Appendix D).
3 SliceFine: Slice as Efficient Fine-tuning
[…]
SliceFine has three design knobs: rank $r$, slice selection, and switching interval $N$. Rank controls capacity; by Corollary 2.6, taking $r\!\geq\!k_{\mathrm{task}}(\tau)$ ensures the active slice intersects the task subspace, yielding a nonzero restricted gradient and a guaranteed decrease (Lemma D.1). We estimate $k_{\mathrm{task}}(\tau)$ once from frozen features on a small calibration set; empirically, a very small rank (often r=1r{=}1) is sufficient for efficient fine-tuning.
For slice selection, spectral balance implies robustness to position: any row or column slice of rank $r$ behaves similarly. We therefore initialize the mask at the first (index-0) position and perform a deterministic sweep across positions, which simplifies implementation and makes the training dynamics easy to track. We also verify that purely random mask placements perform comparably (Appendix K). The interval $N$ trades off per–slice adaptation and coverage: small $N$ explores more positions but can slow early convergence; large $N$ adapts deeper on each slice but delays coverage.
Our ablation E.2 shows a broad, task–dependent sweet spot (e.g., $N\!\approx\!100$$500$ for STS–B/QNLI (Wang et al., 2018) and $N\!\approx\!500$$1500$ for MRPC/SST–2 (Wang et al., 2018)), consistent with the theory that each visited slice contributes additional task–aligned directions until the combined span matches the task dimension. Computationally, per–layer cost scales as $O(d_{\ell}\times r)$ (row) or $O(d_{\ell-1}\times r)$ (column) with zero auxiliary parameters, and optimizer state is maintained only for active entries, yielding favorable memory and throughput characteristics (Details in Section 4). The implementation details are presented in the Appendix L.
4 Experiments
In this section, we conduct experiments across a diverse set of downstream tasks spanning text, image, and video, thereby covering commonsense reasoning, mathematical reasoning, natural language understanding, image classification, and video action recognition. This broad evaluation design allows us to examine whether the theoretical guarantees of the Winner Slice Hypothesis translate into consistent practical gains across modalities. We report results for six variants that differ only in slice rank and orientation: SliceFine-1R, SliceFine-1C, SliceFine-1RC, SliceFine-5R, SliceFine-5C, SliceFine-5RC. Here “1R/5R” denotes a row slice of rank $r\in\{1,5\}$; “1C/5C” denotes a column slice of the same rank; and “1RC/5RC” alternates row and column slices with the same total rank per switch. Additionally, Table 4 reports performance across slice ranks $r\in\{1,2,4,8,32,64,128\}$ (see Appendix E.3 for details) and Figure 3(a) shows increasing higher rank tends to gradually overfit. Across all tables, the best result is highlighted in blue and the second best in orange. The complete experimental setup is described in detail in the appendices. Appendix G provides dataset descriptions, Appendix F lists all baseline PEFT methods, Appendix I outlines hyperparameter choices and training protocols, and Appendix H details the backbone models used in each modality.
Main Results. On commonsense and mathematical reasoning with LLaMA-3B, SliceFine (Table 1) matches or surpasses strong PEFT baselines—including LoRA, AdaLoRA, and RoCoFT—while using fewer trainable parameters. For example, SliceFine-5RC attains the highest average score ($82.13\%$) in math reasoning, outperforming AdaLoRA by +1.07+1.07 points while requiring significantly fewer trainable parameters. Even rank one column or row slice consistently surpasses lighter baselines such as BitFit and Prefix Tuning.
On visual downstream tasks (Table 2), SliceFine matches or outperforms strong low-rank baselines such as VeRA and AdaLoRA, while updating only 0.080.08M–0.410.41M parameters. For instance, on VTAB-1K image classification using ViT-Base-Patch16-224, SliceFine-5R reaches $88.85\%$ average accuracy, surpassing LoRA ($88.08\%$) and AdaLoRA ($87.96\%$). On video recognition using VideoMAE-base, SliceFine-5RC achieves $73.09\%$, outperforming HRA ($72.53\%$) and MiSS ($72.99\%$), highlighting that slice winners generalize beyond static perception to spatio-temporal modeling.
Commonsense Reasoning Math Reasoning LLM Method #TTPs BoolQ PIQA SIQA H.Sw. W.Gra. ARCe ARCc OBQA Avg. M.Ar. G.8K A.S. Se.Eq S.MP Avg. LLaMA-38B Prefix $38.09$ $70.85$ $81.91$ $78.66$ $79.98$ $75.11$ $74.40$ $59.88$ $73.08$ $74.23$ $87.29$ $71.66$ $84.24$ $82.93$ $54.20$ $76.06$ AdaLoRA $33.05$ $71.05$ $82.11$ 82.27 $91.76$ $79.31$ $79.17$ $61.91$ $77.14$ $77.71$ $91.57$ $79.68$ $89.22$ $83.97$ 63.51 $81.41$ VeRA $15.05$ $69.32$ $81.20$ $80.39$ $91.15$ $79.47$ $78.26$ $62.02$ $76.73$ $77.32$ 92.16 $78.36$ $88.91$ $84.48$ $61.71$ $81.07$ LoRA $13.77$ $71.25$ $82.01$ $79.47$ $91.96$ 80.29 79.83 $62.22$ $77.55$ $78.12$ $91.47$ $78.15$ $84.99$ 84.96 $62.02$ $80.32$ RoCoFT $6.90$ 72.47 $82.62$ $80.79$ $91.67$ $79.78$ $79.17$ $62.22$ 79.37 $78.51$ $91.38$ $80.49$ $88.51$ $83.46$ $63.13$ $81.55$ HRA 6.25 $72.27$ 82.82 $80.29$ $91.97$ $79.37$ $79.07$ $62.52$ 79.27 $78.44$ $92.18$ $80.29$ $89.42$ $83.36$ $62.83$ $81.62$ SliceFine-1R 2.18 $70.88$ $81.16$ $79.12$ $90.19$ $78.04$ $78.58$ $61.95$ $78.73$ $77.33$ $91.65$ $79.85$ $88.36$ 84.83 $62.55$ $81.45$ SliceFine-1C 2.18 $70.94$ $81.08$ $78.95$ $90.98$ $78.00$ $77.56$ $61.14$ $77.70$ $77.04$ $91.44$ $78.80$ $87.21$ $82.72$ $61.73$ $80.38$ SliceFine-1RC 2.18 $71.02$ $82.46$ 81.26 92.61 $80.28$ $79.80$ $61.93$ $79.18$ $78.66$ $92.12$ $80.11$ 89.77 $82.18$ $63.55$ $81.55$ SliceFine-5R $6.89$ $71.96$ $82.40$ $81.20$ $92.58$ $80.23$ $79.77$ 62.88 $78.92$ 78.74 92.15 81.05 $89.70$ $84.11$ $63.50$ 82.08 SliceFine-5C $6.89$ 72.49 82.86 $80.97$ $91.97$ $79.71$ 79.82 $62.47$ $78.40$ $78.73$ $92.11$ $80.52$ $89.11$ $83.55$ $63.08$ $81.72$ SliceFine-5RC $6.89$ $72.01$ $82.45$ $81.25$ 92.64 $80.27$ 79.82 62.92 $78.97$ 78.79 $92.10$ 81.10 89.75 $84.16$ 63.53 82.13
Table 1: Commonsense and math reasoning with LLaMA-3B. SliceFine (shaded) rivals or surpasses baselines such as LoRA and AdaLoRA while using far fewer trainable parameters (#TTPs).
Image Video PEFT Caltech Flowers Pets Camel. Euro. Retino. KITTI Avg #TTPs UCF101 Kinetics HMDB Avg #TTPs Full $89.92$ $97.41$ $85.87$ $81.65$ $88.12$ $73.62$ $77.93$ $84.93$ $85.83$ $92.30$ $55.23$ $65.79$ $74.99$ $86.65$ VeRA $91.53$ $99.19$ $91.04$ $86.45$ $92.97$ $74.25$ $77.92$ $87.62$ $0.240$ $92.28$ $57.21$ $66.77$ $72.09$ $0.242$ BitFit $90.58$ $98.93$ $91.06$ $86.73$ $93.07$ $74.39$ $79.26$ $87.72$ 0.101 $92.38$ $57.31$ $66.87$ $72.19$ 0.104 DiffFit $90.26$ $99.24$ 91.73 $86.79$ $92.53$ $74.46$ $81.01$ $88.00$ $0.148$ $92.80$ $57.73$ $67.29$ $72.61$ $0.150$ SHiRA $88.72$ $99.14$ $89.64$ $81.85$ $91.93$ $74.28$ $80.57$ $86.59$ $0.667$ $91.82$ $57.75$ $63.31$ $70.96$ $0.668$ LayerNorm $89.79$ $99.13$ $91.41$ $86.29$ $92.88$ $74.82$ $78.16$ $87.50$ $3.041$ $92.16$ $57.09$ $66.65$ $71.97$ $3.867$ MiSS $89.83$ $99.18$ $91.48$ $86.74$ $91.59$ $73.21$ $85.64$ $88.24$ $0.667$ $93.18$ $58.11$ $67.67$ $72.99$ $0.668$ Propulsion $91.55$ $99.25$ $91.44$ $86.23$ 94.77 $72.36$ $80.27$ $87.96$ $0.165$ $93.47$ $57.40$ $66.96$ $72.61$ $0.166$ LoHA $92.13$ $99.28$ $91.07$ $85.84$ $93.38$ $73.25$ $80.83$ $87.97$ $1.667$ $93.63$ $57.56$ $67.12$ $72.77$ $1.669$ DoRA $91.86$ $99.27$ $91.08$ $85.88$ $91.42$ 75.28 $80.46$ $87.89$ $0.834$ $92.84$ $57.77$ $67.33$ $72.65$ $0.836$ Bone $92.14$ $99.23$ $91.05$ $86.28$ $92.83$ $73.71$ 80.91 $88.02$ $0.414$ $93.82$ $57.75$ $67.31$ $72.96$ $0.415$ RoCoFT $92.56$ $99.31$ $91.19$ 87.84 $93.54$ $74.27$ $79.63$ $88.33$ $0.415$ $93.14$ 58.07 67.63 $72.95$ $0.417$ HRA $92.16$ $99.32$ $91.36$ $86.74$ $93.82$ $74.87$ $78.18$ $88.06$ $0.491$ $92.72$ $57.65$ $67.21$ $72.53$ $0.493$ LoRA $92.03$ $99.18$ $90.92$ 87.73 $92.65$ $74.23$ $80.42$ $88.08$ $0.833$ $93.88$ 57.81 $67.37$ $73.02$ $0.835$ Ada-LoRA $91.62$ $99.24$ $91.18$ $87.54$ $92.37$ $73.84$ $79.94$ $87.96$ $2.011$ 94.43 $57.66$ $67.22$ $73.04$ $2.027$ SliceFine-1R $91.64$ $99.12$ $91.26$ $86.88$ $93.83$ $74.18$ $80.11$ $88.15$ 0.084 $93.14$ $57.16$ $66.29$ $72.20$ 0.087 SliceFine-1C $91.78$ 99.72 $91.12$ $86.63$ $94.24$ $74.10$ $79.73$ $88.19$ 0.084 $93.35$ $57.72$ $66.31$ $72.46$ 0.087 SliceFine-1RC $91.74$ $99.19$ $91.51$ $86.41$ $94.70$ $74.46$ $79.12$ $88.17$ 0.084 $93.36$ $57.10$ $67.30$ $72.59$ 0.087 SliceFine-5R 92.75 99.73 91.64 $87.72$ $94.70$ 75.28 $80.11$ 88.85 $0.415$ $94.54$ $57.16$ 67.49 $73.06$ $0.417$ SliceFine-5C 92.64 $99.64$ $91.15$ $87.78$ $94.42$ $75.14$ 80.39 $88.74$ $0.415$ 94.53 $57.68$ $67.44$ 73.22 $0.417$ SliceFine-5RC $92.28$ $99.53$ $91.71$ $87.58$ 94.89 75.20 $80.89$ 88.83 $0.415$ $94.17$ $57.69$ $67.40$ 73.09 $0.417$
Table 2: Performance comparison on VTAB-1K image and video benchmarks. SliceFine achieves competitive or superior performance to SOTA PEFT baselines with significantly fewer parameters.
Figure 3: Empirical evidence for the robustness of slice selection strategies across tasks. (a) Rank vs. Accuracy: Increasing the slice rank improves accuracy up to a point, after which validation accuracy declines, indicating gradual overfitting. (b) Position vs. Accuracy: accuracy remains stable across slice positions, within $\pm 1\%$ of the anchor accuracy. (c) Wanda category ablations: accuracy is insensitive to whether slices are chosen from most important, less important, mixed, or random weights. (d) LTH comparison: even “bad” slices perform comparably to the “best” slices, supporting the winner-slice property—pretrained networks contain many capable subnetworks.
Is any slice a winner? Empirically, we investigate the Universal Winning–Slice Hypothesis through three complementary experiments, all of which suggest that slice selection is largely insensitive—supporting the winner-slice property. As shown in Figure 3(b), performance remains stable across slice positions. Using a fixed slice rank of $5$, we evaluate accuracy when training slices at different positions of the weight matrix. Accuracy remains within $\pm 1\%$ of the anchor across all positions, indicating that row and column slices contribute comparably to the task-relevant subspace.
[…]
Finally, Figure 3(d) compares “good” and “bad” slices using the Lottery Ticket Hypothesis framework following Frankle & Carbin (2019). Standard LTH seeks a sparse binary mask $M\in\{0,1\}^{d}$ such that the pruned subnetwork $\theta\odot M$ matches the accuracy of the full model: $\mathcal{L}(f_{\theta\odot M}(x),y)\approx\mathcal{L}(f_{\theta}(x),y).$ We extract both “winning” and “losing” subnetworks and use their masks to define slices. Surprisingly, even slices derived from “bad” subnetworks perform comparably to those from “good” ones, further reinforcing that pretrained networks contain many capable subnetworks and that every slice can be a winner.
Figure 4: Comparison of PEFT methods on (a) PEFT model size, (b) peak memory, (c) throughput, and (d) total training time across ViT, VideoMAE, RoBERTa, and multiple datasets.
Methods Space Time #TTPs #APs
FT $O(d_{\ell}\times d_{\ell-1})$ $O(d_{\ell}\times d_{\ell-1})$ $d_{\ell}\cdot d_{\ell-1}$ 0
$(\text{IA})^{3}$ (Liu et al., 2022) $O(d_{k}+d_{v}+d_{ff})$ $O(d_{k}+d_{v}+d_{ff})$ $d_{k}+d_{v}+d_{ff}$ $d_{k}+d_{v}+d_{ff}$
Prompt (Lester et al., 2021) $O(d_{\ell}\times l_{p})$ $O(d_{\ell}\times l_{p})$ $l_{p}\cdot d_{\ell}$ $l_{p}\cdot d_{\ell}$
Prefix (Li & Liang, 2021) $O(L\times d_{\ell}\times l_{p})$ $O(L\times d_{\ell}\times l_{p})$ $L\cdot l_{p}\cdot d_{\ell}$ $L\cdot l_{p}\cdot d_{\ell}$
LoRA (Hu et al., 2021) $O((d_{\ell}+d_{\ell-1})\times r)$ $O((d_{\ell}+d_{\ell-1})\times r)$ $2\cdot d_{\ell}\cdot r$ $2\cdot d_{\ell}\cdot r$
LoRA-FA (Zhang et al., 2023a) $O((d_{\ell}+d_{\ell-1})\times r)$ $O((d_{\ell}+d_{\ell-1})\times r)$ $d\cdot r$ $2\cdot d\cdot r$
AdaLoRA (Zhang et al., 2023b) $O((d_{\ell}+d_{\ell-1}+r)\times r)$ $O((d_{\ell}+d_{\ell-1}+r)\times r)$ $2\cdot d\cdot r+r^{2}$ $2\cdot d\cdot r+r^{2}$
LoHA (Hyeon-Woo et al., 2021) $O(2r\times(d_{\ell}+d_{\ell-1}))$ $O(2r\times(d_{\ell}+d_{\ell-1}))$ $4\cdot d_{\ell}\cdot r$ $4\cdot d_{\ell}\cdot r$
Propulsion (Kowsher et al., 2024b) $O(d)$ $O(d)$ $d$ $d$
SliceFine (row) $O(d_{\ell}\times r)$ $O(d_{\ell}\times r)$ $r\cdot d_{\ell}$ 0
SliceFine (column) $O(d_{\ell-1}\times r)$ $O(d_{\ell-1}\times r)$ $r\cdot d_{\ell-1}$ 0
Table 3: Comparison of space and time complexity, total trainable parameters (#TTPs), and additional parameters (#APs) introduced by different PEFT methods for a single layer $W^{(\ell)}\in\mathbb{R}^{d_{\ell}\times d_{\ell-1}}$. We denote $d_{k},d_{v},$ and $d_{ff}$ as the dimensions of the three learned vectors in IA3, and $l_{p}$ as the prompt length in prompt tuning and prefix tuning. For LoRA-type methods, $r$ denotes the rank. SliceFine achieves $O(d_{\ell}\times r)$ or $O(d_{\ell-1}\times r)$ complexity with no additional parameters, in contrast to other methods that incur higher asymptotic costs or parameter overhead.
Efficiency Analysis. In this section, we discuss the efficiency of SliceFine. All methods are trained under identical conditions: 10 epochs, batch size 64, fixed sequence lengths, BF16 precision, and the same learning schedules. Reported runtime is the wall-clock time for a full training run, averaged over three seeds, and throughput is measured in iterations per second on identical hardware (NVIDIA A100 80GB). Table 3 summarizes asymptotic costs. Slice training scales as $O(d_{\ell}\times r)$ (row) or $O(d_{\ell-1}\times r)$ (column) and introduces no additional parameters (#APs = 0). In contrast, LoRA-style and adapter-style baselines incur $2r(d_{\ell}+d_{\ell-1})$ or higher parameter overhead, significantly increasing both space and time complexity.
[…]
Overall: (i) SliceFine supports the UWSH means any slice with sufficient rank acts as a local winner, so training slice suffices for PEFT without adding adapters; (ii) empirically, a rank–1 slice already delivers competitive performance; and (iii) SliceFine achieves comparable or superior accuracy to the baseline while requiring significantly fewer trainable parameters, resulting in faster training and lower memory consumption. More detailed results are provided in Appendix J. In addition, we present detailed ablation studies in Appendix E, including the effect of slice rank (E.1) and the switching interval (E.2).
5 Related Work
Lottery Ticket Hypothesis. LTH (Frankle & Carbin, 2019) says a large dense network contains sparse subnetworks (“winning tickets”) that, after pruning and retraining, can match the full model. Later work tested this on modern architectures (Burkholz, 2022) and surveyed the evidence (Liu et al., 2024a). Pruning methods to find these subnetworks were studied in Zhou et al. (2019). The strong LTH (Ramanujan et al., 2020) shows that some untrained sparse subnetworks already perform well, with further analysis and CNN results in Pensia et al. (2020); da Cunha et al. (2022). In pretrained models, similar findings appear: BERT has winning-ticket subnetworks (Chen et al., 2020; Yu et al., 2021); pruning before fine-tuning can find them (Wu et al., 2022); and ImageNet-pretrained CNNs also contain sparse subnetworks that keep downstream accuracy (Chen et al., 2021; Iofinova et al., 2022).
PEFT. PEFT adapts large models by updating a small subset of parameters while freezing most of the backbone. Prior work spans prompt-based approaches that optimize soft prompts or prefixes (Lester et al., 2021; Li & Liang, 2021), adapter layers inserted into residual blocks (Houlsby et al., 2019), and weight-space updates such as LoRA (Hu et al., 2021), BitFit and LayerNorm tuning (Zaken et al., 2021; Zhao et al., 2023), and multiplicative gating like IA3 (Liu et al., 2022). Numerous refinements improve placement, rank, and overhead: AdaLoRA adapts rank during training (Zhang et al., 2023b), LoRA-FA and VeRA reduce parameter and optimizer cost (Zhang et al., 2023a; Kopiczko et al., 2023), and other variants explore alternative factorizations or injection points (Hyeon-Woo et al., 2021; Edalati et al., 2022; Bałazy et al., 2024). Beyond LoRA, DoRA stabilizes updates via weight decomposition (Liu et al., 2024b), READ augments with residual adapters (Wang et al., 2024), LoRA+ and KronA target better efficiency (Hayou et al., 2024; Edalati et al., 2022), and Q-PEFT conditions adaptation on task queries (Peng et al., 2024). Orthogonal lines combine PEFT with sparsity: structured and unstructured pruning (e.g., SparseGPT, Wanda, LLM-Pruner) remove redundancy with limited loss (Frantar & Alistarh, 2023; Sun et al., 2023a; Ma et al., 2023), and domain adaptation strategies help avoid overfitting (Gururangan et al., 2020). Closer to subnetwork training, RoCoFT trains a fixed subset of rows/columns and reports strong results, but the chosen subnetwork remains static over training (Kowsher et al., 2024a).
In contrast to methods that add auxiliary parameters (adapters, prompts, low-rank factors) or remove weights via pruning, SliceFine updates a set of slices across layers of the original network—without adding parameters—and can move the active slice during training, yielding block–coordinate adaptation motivated by spectral balance and high task energy in the pretrained backbone.
6 Conclusion
This work proposes the Universal Winning–Slice Hypothesis (UWSH): in pretrained networks satisfying spectral balance and high task energy, every sufficiently wide slice of a weight matrix acts as a local winning ticket, and a small set of slices across layers forms a global winning ticket. Building on this view, SliceFine fine-tunes only a small, moving slice per layer—adding no new parameters—while matching strong PEFT baselines in accuracy and improving efficiency in memory, speed, and model size. Experiments across text, image, and video tasks show competitive performance compared to SOTA PEFT methods.
Appendix A Proof of Theorem Universal Winning Ticket
Proof of Theorem 2.4.
Let $\mathcal{L}(\theta)=\tfrac{1}{n}\sum_{i=1}^{n}\ell(f_{\theta}(x_{i}),y_{i})$ denote the empirical downstream loss on dataset $D=\{(x_{i},y_{i})\}_{i=1}^{n}$, with pretrained initialization $\theta_{0}$. For any slice $M\subseteq W^{(\ell)}$, define the restricted gradient
$g_{M}\;=\;\nabla_{M}\mathcal{L}(\theta)=\tfrac{\partial\mathcal{L}(\theta)}{\partial(M\odot W^{(\ell)})}=\tfrac{1}{n}\sum_{i=1}^{n}J_{M}(x_{i})^{\top}\nabla_{f}\ell(f_{\theta}(x_{i}),y_{i}),$ (1)
where $J_{M}(x)=\tfrac{\partial f_{\theta}(x)}{\partial(M\odot W^{(\ell)})}\in\mathbb{R}^{k\times|M|}$ is the slice Jacobian and $\nabla_{f}\ell(\cdot,y_{i})\in\mathbb{R}^{k}$ is the gradient with respect to logits. By assumption there exists at least one $i$ with $\nabla_{f}\ell(f_{\theta_{0}}(x_{i}),y_{i})\neq 0$, and by the projection condition each $J_{M}(x)$ has nontrivial overlap with the $k_{\mathrm{task}}$-dimensional task subspace. Hence $g_{M}\neq 0$ for all nonempty $M$.
Assume slice-restricted Lipschitz smoothness: there exists $L_{M}>0$ such that
$\mathcal{L}(\theta_{0}+M\odot\Delta)\leq\mathcal{L}(\theta_{0})+g_{M}^{\top}\mathrm{vec}(\Delta)+\tfrac{L_{M}}{2}\|\mathrm{vec}(\Delta)\|_{2}^{2}$ (2)
for all $\Delta$ with $\|\Delta\|_{2}\leq\rho$. Choosing $\Delta=-\alpha g_{M}$ yields
$\mathcal{L}(\theta_{0}-\alpha M\odot g_{M})\leq\mathcal{L}(\theta_{0})-\alpha\|g_{M}\|_{2}^{2}+\tfrac{L_{M}}{2}\alpha^{2}\|g_{M}\|_{2}^{2},$